Problem 2-2

The manager of a crew that installs carpeting has tracked the crew’s output over the past several weeks, obtaining these figures:

   

Week Crew Size Yards Installed
1 4 97
2 3 69
3 4 99
4 2 52
5 3 63
6 2 52

    

a. Compute the labor productivity for each of the weeks. (Round your answers to 2 decimal places.)

    

Week Crew size     Labor productivity
   (Yards/Person)
1 4     
2 3     
3 4     
4 2     
5 3     
6 2     


   

b. Which crew size works best?

    

A crew of    (Click to select)  3  4  2   seems to work best.
Answer #1

PARTIAL PRODUCTIVITY = TOTAL OUTPUT / PARTICULAR INPUT


WEEK

CREW SIZE

INSTALLATION

PRODUCTIVITY

FORMULA

1

4

97

24.25

97 / 4 = 24.25

2

3

69

23

69 / 3 = 23

3

4

99

24.75

99 / 4 = 24.75

4

2

52

26

52 / 2 = 26

5

3

63

21

63 / 3 = 21

6

2

52

26

52 / 2 = 26

BASED ON THE ABOVE TABLE, THE SIZE OF 2 CREW MEMBERS IS THE MOST PRODUCTIVE.